Consider the following data set:
10 8 9 13 7
If two numbers are randomly selected, what is the probability that their average will be less than 9.75?

Answer by stanbon(75887)

Consider the following data set:
10 8 9 13 7
If two numbers are randomly selected, what is the probability that their average will be less than 9.75?
----
10;8... ave = 18/2 = 9
10;9... ave =........9.5
10:13.. ave = .......11.5
10:7... ave = ....... 8.5
8:9.... ave = ....... 8.5
8:13... ave = .......10.5
8:7.... ave = ....... 7.5
9;13... ave = ........11
9:7.... ave = ........ 8
13:7... ave = ........10
----
P(average < 9.75) = 6/10
===============================
Cheers,
Stan H.
========================